# TongUI: Building Generalized GUI Agents by Learning from Multimodal Web Tutorials

### v1

@@ Abstract
Building Graphical User Interface (GUI) agents is a promising research direction, which simulates human interaction with computers or mobile phones to perform diverse GUI tasks. However, a major challenge in developing generalized GUI agents is the lack of sufficient trajectory data across various operating systems and applications, mainly due to the high cost of manual annotations. In this paper, we propose the TongUI framework that builds generalized GUI agents by learning from rich multimodal web tutorials. Concretely, we crawl and process online GUI tutorials (such as videos and articles) into GUI agent trajectory data, through which we produce the GUI-Net dataset containing 143K trajectory data across five operating systems and more than 200 applications. We develop the TongUI agent by fine-tuning Qwen2.5-VL-3B/7B models on GUI-Net, which show remarkable performance improvements on commonly used grounding and navigation benchmarks, outperforming baseline agents about 10% on multiple benchmarks, showing the effectiveness of the GUI-Net dataset and underscoring the significance of our TongUI framework. We will fully open-source the code, the GUI-Net dataset, and the trained models soon.
Figure 1: TongUI crawls and processes multimodal web tutorials from diverse sources into GUI trajectories. The collected dataset spans over 200 applications across five operating systems, leading to powerful generalization in a wide range of GUI tasks.
footnotetext: * Equal contribution.    🖂  Corresponding author.
@@ 1 Introduction
Graphical User Interface (GUI) agents based on large foundation models are designed to automate tasks on digital devices by emulating human interactions with a variety of applications on different devices [23, 36, 44]. These agents utilize large language models (LLMs) or vision-language models (VLMs) to process visual inputs (screenshots) and textual inputs (accessibility tree and HTML code) and produce corresponding actions, such as clicking buttons, filling forms, and scrolling, to complete GUI tasks that are often repetitive and time-consuming [14, 10, 43]. GUI agents significantly enhance human-computer interaction, providing potential applications to various domains, such as software testing, financial services, office assistance, and industrial automation, improving work efficiency and user experience.
Recently, notable efforts have been made for GUI agents by fine-tuning the foundation models [17, 27, 40]. However, collecting sufficient data for fine-tuning is challenging. Most existing approaches rely on either manually annotated interaction trajectories [19], which are high-quality but costly to obtain, or synthetic data generated from large open-source or proprietary LLM/VLMs, which may lack diversity and accuracy [26, 20, 6]. Desirable GUI agents need to perform actions across a wide variety of applications, each with unique interaction sequences, but collecting comprehensive and diverse data across different applications and operating systems remains a major challenge. As a result, the lack of large-scale, diverse, and well-structured GUI data continues to be a key bottleneck in developing robust and generalized GUI agents.
@@
In doing so, we design the TongUI framework composed of three steps: tutorial crawling, tutorial processing, and trajectory generation. In the tutorial crawling step, we write some seed tasks and use LLMs to extend them into a wider collection of tasks, ensuring comprehensive coverage and enhanced realism of real-world GUI tasks. The generated tasks serve as keywords for retrieving content from hosts for online GUI tutorials (such as articles from WikiHow and videos from YouTube). The tutorial processing step aims to extract the textual descriptions and screenshots of multimodal tutorials. We first obtain textual descriptions of multimodal tutorials via automatic speech recognition (ASR) or captioning, through which task queries and plans are produced using LLMs on the obtained textual descriptions. Then, we extract salient frames from videos as the screenshot of each step, while the images in articles are directly regarded as screenshots. Finally, we leverage one zero-shot GUI agent to automatically recognize trajectories, including reasoning thoughts and actions between two steps. Based on the tutorial collection pipeline, we construct a GUI-Net dataset that contains 143K data across five operating systems with more than 200 applications.
Based on GUI-Net, we develop the TongUI agent using Qwen2.5-VL-3B/7B models. We evaluate the TongUI agents on both the offline and online settings, and the results show that the TongUI agent exhibits consistent improvements in grounding and navigation capabilities. The results demonstrate the generalization of the TongUI agent, since it is trained on high-quality data across diverse applications, underscoring the effectiveness of the TongUI framework that improves the agents in a low-cost manner without the need for expensive manual annotation.
@@
  • We propose the TongUI framework that enables GUI agents to automatically learn from web resources via tutorial crawling, tutorial processing, trajectory generation, and agent tuning, leading to better generalization.
  • We construct the GUI-Net dataset, containing 143K data across five operating systems and over 200 applications.
  • We develop the TongUI agent by fine-tuning Qwen2.5-VL models on GUI-Net, achieving improvements on multiple popular benchmarks.
We will fully open-source the code, the GUI-Net dataset, and the trained models.
@@ 2.2 Agent Data Collection
Compared with data collection for LLMs or VLMs, collecting high-quality data for agents is more challenging, since agents usually require long trajectories that solve complex tasks in different domains (e.g., GUI, multimodal reasoning, embodied AI). Commonly used schemes include human annotation and model synthesis. Human annotation could produce high-quality and diverse trajectory data, but may result in high cost [19]. Some efforts have been made to use zero-shot agents to produce synthetic trajectory data [26, 20, 6], while this may cause out-of-distribution data, leading to inferior results. Similar to other domains, human annotation and model generation are also commonly used to collect trajectories for GUI agents, including the planning, action sequence, and action position [5, 31, 3, 27, 4, 21]. In addition to the two schemes, considering that there are abundant resources for GUI operations on the Internet, some efforts focus on collecting trajectory data of GUI agents from the Web, to preserve the data quality and reduce costs simultaneously [24, 39]. They collect textual tutorials from the Internet, and convert them into trajectories by using LLMs or exploring in simulators. Unlike them, we directly collect multimodal tutorials from the Internet and propose a multimodal tutorial process pipeline to convert them into trajectories without any simulators. Meanwhile, compared with synthetic data in simulators, our data is closer to tasks in the real world with practical purposes, benefiting to the generalization capability of GUI agents.
@@ 3 Method
Figure 2: Illustration for TongUI model architecture.
Figure 3: Illustration of TongUI, including tutorial source, tutorial crawling, tutorial process, trajectory generation, and agent tuning.
@@ 3.1 Formulation
We formulate GUI tasks as a sequential decision process, and we adopt ReAct [42] as our agent framework. At one time step $i$, given the observation of previous $n$ steps $o_{i-n},\cdots,o_{i-1},o_{i}$ (e.g., a screenshot of GUI), the thoughts $r$ and actions $a$ of previous $n$ steps $(r_{i-n},a_{i-n},\cdots,r_{i-1},a_{i-1})$, and the query $q$, the agent generates a new thought $r_{i}$ and an executable action $a_{i}$ from the action space, such as clicking on a specific UI element, entering text, or scrolling through the interface. Then, executing the action $a_{i}$ leads to a new observation $o_{i+1}$ (e.g., an updated screenshot). This interaction loop continues, with the agent repeatedly observing the environment, selecting actions, and receiving updated observations until either a termination condition is met (e.g., the task is completed or fails) or the maximum number of steps is reached. We parameterize the GUI agent using a vision-language-action model $M_{\theta}$,
$r_{i}^{\star},a_{i}^{\star}=\arg\max_{r_{i},a_{i}}M_{\theta}(r_{i},a_{i}|$ $q,o_{i-n},r_{i-n},a_{i-n}\cdots,$ (1)
$o_{i-1},r_{i-1},a_{i-1},o_{i}).$
The architecture of the used model is shown in Fig. 2.
@@ 3.3 Tutorial Crawling
We use a keyword-based search approach. The search keywords are constructed by combining the name of the applications or website with the task content, that is “app/web + task.” Here, “app/web” refers to the operating system or application being used (e.g., a mobile app or website), while “task” refers to a complete workflow or a set of actions required to accomplish a specific objective (e.g., changing font size in Word, or browsering in Chrome). We identify some task seeds across multiple applications via brainstorming, and expanded these seeds using LLMs to generate more diverse and relevant tasks.
@@ 3.4 Tutorial Process
Textual processing. Given a crawled multimodal tutorial $\{v,e\}$, this process extracts the task $q$ and rough descriptions $\{h_{1},\cdots,h_{T}\}$ in $T$ steps. In video tutorials from Youtube, Tiktok, and Bilibili, the task guidance is usually in the video’s audio information without any textual introductions. To solve this issue, we begin the processing by applying the open-source speech recognition model, Whisper [28], to transcribe the audio streams into text. This step enables us to extract the text information $e$ from these video tutorials. For image-text articles on Wikihow and Baidu Experience, we develop specific parsing tools for each website based on their article structures to extract the textual information $e$.
We use LLMs to clean the text $e$ by removing irrelevant information, such as advertisements, off-topic content, and redundant expressions. LLMs are also used to identify key task-related verbs and nouns, such as “clicking a button”, “filling out a form”, or “selecting a menu,” which are crucial for understanding the tasks involved in the tutorial. Then, we leverage both the title and extracted text to classify tutorials into three categories: mobile, desktop, and other. The category ‘other’ indicates this task is irrelevant to GUI tasks. The categories ‘mobile’ and ‘desktop’ allow us to organize the data by device types, ensuring the tasks are appropriately tailored to their respective environments. Finally, we use LLMs to extract the task $q$ and summarize this process into rough descriptions $\{h_{1},\cdots,h_{T}\}$ of $T$ steps from the filtered textual information.
@@
For video tutorials, we need to extract key frames that correspond to critical actions in video tutorials, representing meaningful moments in the task-solving process. This process is crucial for constructing training data from video tutorials. We observe that the audio transcript of the videos often contains valuable planning information, such as the description of task steps. Therefore, when audio is available, we first segment the video based on the audio transcript timestamps, treating each segment as a potential task step. We then extract key frames from each of these segments. In the absence of audio, we treat the entire video as a single segment and extract key frames throughout the video. In doing so, we adopt the MOG2 background subtraction algorithm [48, 49, 50] to detect significant changes in GUI interactions. Specifically, this algorithm models the static background of GUI and isolates dynamic foreground elements that correspond to user interactions, such as clicks, form submissions, or UI transitions. This algorithm updates the background model continuously as it processes consecutive video frames, detecting changes between each frame and the background. When significant differences between the current frame and the background are detected, the frame is considered a key frame, capturing potential GUI action. The frames preceding these detected changes are then considered as representative observations of the interface before the critical action occurs.
Figure 4: Data statistics on the GUI-Net dataset.
@@ 3.5 Trajectory Generation
This process aims to generate the trajectory $(o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T})$ in $T$ steps using the task $q$, and image $o_{i}$ and rough description $h_{i}$ in $T$ steps. Concretely, we use a pretrained GUI agent (such as UI-TARS [27]) to generate the thought $r_{i}$ and the action $a_{i}$ using the same action space defined by the GUI agent. For the thought $r_{i}$ and action $a_{i}$, we feed the observation $o_{i}$ with $h_{i}$ as the query to the zero-shot agent. Here, we use $h_{i}$ as the query for agents instead of the task $q$, because $q$ usually contains an abstract goal instead of specific instruction about what to do on the screen. We empirically find that using $h_{i}$ leads to better performance. In this case, each step of text and image pairs can be written as $\{o_{i},r_{i},a_{i}\}$. In practice, the model might fail to generate well-formatted actions. We discard this step and make the next action the beginning of a new trajectory. For example, for a 4-step trajectory which has a failed generation in the second step, we split it into two training trajectories: ($(\{o_{1},r_{1},a_{1}\})$ and ($(\{o_{3},r_{3},a_{3}\},\{o_{4},r_{4},a_{4}\})$. Finally, we combine the trajectory with the task $q$ as $(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})$. We collect these trajectories into a GUI-Net dataset that contains 143K data with 1.24B tokens. These data is on five operating systems with more than 200 applications. After that, we use GUI-Net for supervised fine-tuning (SFT).
@@ 3.6 Agent Tuning
$\min\mathbb{E}_{(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})\sim\mathbb{D}}\big{[}-$ $\sum_{i=1}^{T}M_{\theta}(r_{i},a_{i}|q,o_{i-n},r_{i-n},$
$a_{i-n},\cdots,o_{i-1},r_{i-1},a_{i-1},o_{i})\big{]},$
where $\mathbb{D}$ is the collected GUI-Net dataset, and we sum the loss values of the $T$ steps in the trajectory. After training, we obtain the TongUI agent.
@@ 4 Dataset Statistics
We provide four key statistics: operating system distribution, application distribution, step distribution, and action distribution to show the diversity of the collected GUI-Net dataset. We show the operating system distribution in  Fig. 4(a). Our dataset covers a diverse range of operating systems, including Windows, Android, iOS, MacOS, and Linux, ensuring a broad representation of GUI interactions across both desktop and mobile environments. The application distribution is shown in  Fig. 4(b). Our dataset covers a wide variety of application categories, ensuring a broad and diverse range of user tasks. We show the step distribution in  Fig. 4(c). Our dataset includes GUI interactions with varying step lengths, ranging from single-step actions to 9-step tasks. The distribution shows that shorter tasks (1-3 steps) are more frequent, while longer tasks gradually decrease in numbers. This distribution reflects a balance between simple and complex GUI interactions, capturing both quick actions and more involved workflows. The action distribution is shown in  Fig. 4(d). Our dataset encompasses a diverse range of GUI actions, with Click and Tap being the most frequent, reflecting their central role in both desktop and mobile interactions. Other common actions include HotKey, Swipe, and Drag, covering various forms of user interactions.
Model Data Num Data Size Desktop Icon Desktop Text Mobile Icon Mobile Text Web Icon Web Text Average
SeeClick-9.6B [4] 364K - 30.0 72.2 52.0 78.0 32.5 55.7 53.4
UGround-7B [7] 1.3M - 63.6 82.5 60.3 82.8 80.4 73.3 70.4
OmniParser-GPT-4V [22] - - 63.6 91.3 57.0 93.9 51.0 81.3 73.0
ShowUI-2B [17] 256K 0.72B 61.1 76.3 75.5 92.3 63.6 81.7 75.1
UI-TARS-2B [27] - 50B 68.6 90.7 75.5 93.0 74.8 84.3 82.3
UI-TARS-7B [27] - 50B 85.7 95.9 85.2 94.5 83.5 90.0 89.5
Qwen2.5-VL-3B $\dagger$ - - 7.8 22.2 5.2 8.4 1.7 2.4 8.0
Qwen2.5-VL-7B $\dagger$ - - 16.4 26.8 5.2 6.6 7.3 13.0 12.6
TongUI-3B 399K 1.24B 68.5 86.5 76.0 90.5 68.4 87.4 79.6
TongUI-7B 399K 1.24B 75.0 91.2 79.9 93.0 72.3 88.7 83.4
@@
Table 1: Results on ScreenSpot; $\dagger$ means the results are re-produced. We report results on six splits of ScreenSpot and the average scores. The best method is marked in bold, and the second-best method is underlined.

### v2

@@ Abstract
Building Graphical User Interface (GUI) agents is a promising research direction, which simulates human interaction with computers or mobile phones to perform diverse GUI tasks. However, a major challenge in developing generalized GUI agents is the lack of sufficient trajectory data across various operating systems and applications, mainly due to the high cost of manual annotations. In this paper, we propose the TongUI framework that builds generalized GUI agents by learning from rich multimodal web tutorials. Concretely, we crawl and process online GUI tutorials (such as videos and articles) into GUI agent trajectory data, through which we produce the GUI-Net-1M dataset containing 1M trajectory data across five operating systems and more than 280 applications. We develop the TongUI agent by fine-tuning Qwen2.5-VL-3B/7B models on GUI-Net-1M, which shows remarkable performance improvements on commonly used grounding and navigation benchmarks, outperforming baseline agents about 10% on multiple benchmarks, showing the effectiveness of the GUI-Net-1M dataset and underscoring the significance of our TongUI framework. We have fully opened-source the code, the GUI-Net-1M dataset, and the trained models.
Figure 1: TongUI crawls and processes multimodal web tutorials from diverse sources into GUI trajectories. The collected GUI-Net-1M dataset spans over 280 applications across five operating systems, leading to powerful generalization in a wide range of GUI tasks.
@@ 1 Introduction
Graphical User Interface (GUI) agents based on large foundation models are designed to automate tasks on digital devices by emulating human interactions with a variety of applications on different devices [24, 37, 46]. These agents utilize large language models (LLMs) or vision-language models (VLMs) to process visual inputs (screenshots) and textual inputs (accessibility tree and HTML code) and produce corresponding actions, such as clicking buttons, filling forms, and scrolling, to complete GUI tasks [14, 10, 45]. GUI agents significantly enhance human-computer interaction, providing potential applications to various domains, such as software testing, financial services, office assistance, and industrial automation, improving work efficiency and user experience.
Recently, notable efforts have been made for GUI agents by fine-tuning the foundation models [18, 28, 42]. However, collecting sufficient data for fine-tuning is challenging. Most existing approaches rely on either manually annotated interaction trajectories [20] that are high-quality but costly to obtain, or synthetic data generated from large open-source or proprietary LLM/VLMs, which may lack diversity and accuracy [27, 21, 6]. Desirable GUI agents need to perform actions across a wide variety of applications, each with unique interaction sequences, but collecting comprehensive and diverse data across different applications and operating systems remains a major challenge. As a result, the lack of large-scale, diverse, and well-structured GUI data continues to be a key bottleneck in developing robust and generalized GUI agents.
@@
In doing so, we design the TongUI framework composed of four steps: tutorial crawling, tutorial processing, trajectory generation, and data filtering. In the tutorial crawling step, we write some seed tasks and use LLMs to extend them into a wider collection of tasks. The generated tasks serve as keywords for retrieving content from hosts for online GUI tutorials (such as articles from WikiHow and videos from YouTube). The tutorial processing step aims to extract the textual descriptions and screenshots of multimodal tutorials. We first obtain textual descriptions of multimodal tutorials via automatic speech recognition (ASR) or captioning, through which task queries and plans are produced using LLMs on the obtained textual descriptions. Then, we extract salient frames from videos as screenshots of each step, while the images in articles are directly regarded as screenshots. Finally, we leverage one zero-shot GUI agent to automatically recognize trajectories, including reasoning thoughts and actions between two steps. In the data filtering step, we apply a multi-stage pipeline to ensure data quality and GUI relevance, including duplicate tutorial removal, LLM-based content filtering, and trajectory-level filtering using GUI agents. Based on the tutorial collection pipeline, we construct a GUI-Net-1M dataset that contains 1M data across five operating systems with more than 280 applications. Notably, our dataset includes tutorials from different time periods, where evolving layouts still preserve consistent execution logic between similar tasks, helping the model generalize to diverse GUI environments.
Based on GUI-Net-1M, we develop the TongUI agent using Qwen2.5-VL-3B/7B models. We evaluate the TongUI agents on both the offline and online settings, and the results show that the TongUI agent exhibits consistent improvements in grounding and navigation capabilities. The results demonstrate underscores the effectiveness of the TongUI framework and the collected GUI-Net-1M dataset that improve the agents in a low-cost manner without the need for expensive manual annotation.
@@
(1) We propose the TongUI framework that enables GUI agents to automatically learn from web resources via tutorial crawling, tutorial processing, trajectory generation, data filtering, and agent tuning, leading to better generalization.
(2) We construct the GUI-Net-1M dataset, containing 1M data across five operating systems and over 280 applications. We have fully opened-source the code, the GUI-Net-1M dataset, and the trained models.
(3) We develop the TongUI agent by fine-tuning Qwen2.5-VL models on GUI-Net-1M, achieving improvements on multiple popular benchmarks.
@@ 2.2 Agent Data Collection
Compared with data collection for LLMs or VLMs, collecting high-quality data for agents is more challenging, since agents usually require long trajectories that solve complex tasks in different domains (e.g., GUI, multimodal reasoning, embodied AI). Commonly used schemes include human annotation and model synthesis. Human annotation could produce high-quality and diverse trajectory data, but may result in high cost [20]. Some efforts have been made to use zero-shot agents to produce synthetic trajectory data [27, 21, 6], while this may cause out-of-distribution data, leading to inferior results. Similar to other domains, human annotation and model generation are also commonly used to collect trajectories for GUI agents, including the planning, action sequence, and action position [5, 32, 3, 28, 4, 22]. Note that the recent state-of-the-art method UI-TARS [28] has not release its data. In contrast, we will fully release our data.
In addition to the two schemes, considering that there are abundant resources for GUI operations on the Internet, some efforts focus on collecting trajectory data of GUI agents from the Web, to preserve the data quality and reduce costs simultaneously [25, 41]. They collect textual tutorials from the Internet, and convert them into trajectories by using LLMs or exploring in simulators. Unlike them, we directly collect multimodal tutorials from the Internet and propose a multimodal tutorial process pipeline to convert them into trajectories without any simulators. Meanwhile, compared with synthetic data in simulators, our data is closer to tasks in the real world with practical purposes, benefiting to the generalization capability of GUI agents.
Figure 2: Illustration for TongUI model architecture.
@@ 3.1 Formulation
We formulate GUI tasks as a sequential decision process, and we adopt ReAct [44] as our agent framework. At one time step $i$, given the observation of previous $n$ steps $o_{i-n},\cdots,o_{i-1},o_{i}$ (e.g., a screenshot of GUI), the thoughts $r$ and actions $a$ of previous $n$ steps $(r_{i-n},a_{i-n},\cdots,r_{i-1},a_{i-1})$, and the query $q$, the agent generates a new thought $r_{i}$ and an executable action $a_{i}$ from the action space, such as clicking on a specific UI element, entering text, or scrolling through the interface. Then, executing the action $a_{i}$ leads to a new observation $o_{i+1}$ (e.g., an updated screenshot). This interaction loop continues, with the agent repeatedly observing the environment, selecting actions, and receiving updated observations until either a termination condition is met (e.g., the task is completed or fails) or the maximum number of steps is reached. We parameterize the GUI agent using a vision-language-action model $M_{\theta}$ as shown in Fig. 2,
$r_{i}^{\star},a_{i}^{\star}=\arg\max_{r_{i},a_{i}}M_{\theta}(r_{i},a_{i}|q,o_{i-n},r_{i-n},a_{i-n}\cdots,o_{i-1},r_{i-1},a_{i-1},o_{i}).$ (1)
@@ 3.3 Tutorial Crawling
We use a keyword-based search approach. The search keywords are constructed by combining the name of the applications or website with the task content, that is “app/web + task.” Here, “app/web” refers to the operating system or application being used (e.g., a mobile app or website), while “task” refers to a complete workflow or a set of actions required to accomplish a specific objective (e.g., changing font size in Word, or browsering in Chrome). We identify some task seeds across multiple applications via brainstorming, and expanded them using LLMs for more diverse and relevant tasks.
@@ 3.4 Tutorial Process
Textual processing. Given a crawled multimodal tutorial $\{v,e\}$, this process extracts the task $q$ and rough descriptions $\{h_{1},\cdots,h_{T}\}$ in $T$ steps. In video tutorials from Youtube, Tiktok, and Bilibili, the task guidance is usually in the video’s audio information without any textual introductions. To solve this issue, we begin by applying the open-source speech recognition model, Whisper [29], to transcribe the audio streams into the text information $e$. For image-text articles on Wikihow and Baidu Experience, we parse websites based on their article structures for the textual information $e$.
We use LLMs to identify key task-related verbs and nouns, such as “clicking a button”, “filling out a form”, or “selecting a menu”. Then, we leverage both the title and extracted text to classify tutorials into three categories: mobile, desktop, and others. The data with category “others” is discarded. The categories “mobile” and “desktop” allow us to organize the data by device types, ensuring the tasks are appropriately tailored to their respective environments. Finally, we use LLMs to extract the task $q$ and summarize this process into rough descriptions $\{h_{1},\cdots,h_{T}\}$ of $T$ steps from the filtered textual information.
@@
For video tutorials, we need to extract key frames that correspond to critical actions, representing meaningful moments in the task-solving process. We observe that the audio transcript of the videos often contains valuable planning information, such as the description of task steps. Therefore, when audio is available, we first segment the video based on the audio transcript timestamps, treating each segment as a potential task step. We then extract key frames from each of these segments. In the absence of audio, we treat the entire video as a single segment and extract key frames throughout the video. In doing so, we adopt the MOG2 background subtraction algorithm [50, 51, 52] to detect significant changes in GUI interactions.
Figure 4: Data statistics on the GUI-Net-1M dataset.
@@ 3.5 Trajectory Generation
This process aims to generate the trajectory $(o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T})$ in $T$ steps using the task $q$, and image $o_{i}$ and rough description $h_{i}$ in $T$ steps. Concretely, we use a pretrained GUI agent (such as UI-TARS [28]) to generate the thought $r_{i}$ and the action $a_{i}$ using the same action space defined by the GUI agent. For the thought $r_{i}$ and action $a_{i}$, we feed the observation $o_{i}$ with $h_{i}$ as the query to the zero-shot agent. Here, we use $h_{i}$ as the query for agents instead of the task $q$, because $q$ usually contains an abstract goal instead of specific instruction about what to do on the screen. We empirically find that using $h_{i}$ leads to better performance. In this case, each step of text and image pairs can be written as $\{o_{i},r_{i},a_{i}\}$. In practice, the model might fail to generate well-formatted actions. We discard this step and make the next action the beginning of a new trajectory. For example, for a 4-step trajectory which has a failed generation in the second step, we split it into two training trajectories: ($(\{o_{1},r_{1},a_{1}\})$ and ($(\{o_{3},r_{3},a_{3}\},\{o_{4},r_{4},a_{4}\})$. Finally, we combine the trajectory with the task $q$ as $(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})$. We collect these trajectories into a GUI-Net-1M dataset that contains 1M data. These data is on five operating systems with more than 280 applications. After that, we use GUI-Net-1M for supervised fine-tuning (SFT).
3.6 Data Filtering
We apply a multi-stage data filtering pipeline for data quality, as shown in Fig. 3. (1) Duplicate Tutorial Removal. To eliminate redundancy of crawling, we directly remove exact duplicate tutorials based on their unique identifiers (e.g., video IDs and URLs). This step ensures that each retained tutorial is unique, avoiding repetition and simplifying downstream processing. (2) Content Filtering. After tutorial processing, we employ an LLM to semantically filter out tutorials that are irrelevant to the GUI tasks. Concretely, based on the tutorial titles and audio transcriptions, the LLM evaluates whether the content is centered around GUI interactions. (3) Trajectory Filtering. In the trajectory generation stage, since we use UI-TARS [28], if a screenshot corresponds to a step unrelated to GUI interaction, the agent typically predicts actions like wait or call_user. We use this as a signal to discard such observations. Then, we feed the screenshot and trajectory of a task into a Qwen2.5-VL-7B model, and prompt it to filter out low-quality data.
@@ 3.7 Agent Tuning
$\min\mathbb{E}_{(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})\sim\mathbb{D}}\big{[}-\sum_{i=1}^{T}M_{\theta}(r_{i},a_{i}|q,o_{i-n},r_{i-n},a_{i-n},\cdots,o_{i-1},r_{i-1},a_{i-1},o_{i})\big{]},$
where $\mathbb{D}$ is the collected GUI-Net-1M dataset, and we sum the loss values of the $T$ steps in the trajectory. After training, we obtain the TongUI agent.
@@ 4 Dataset Statistics
We provide four key statistics to show the diversity of the collected GUI-Net-1M dataset. We show the operating system distribution in  Fig. 4(a). Our dataset covers a diverse range of operating systems, including Windows, Android, iOS, MacOS, and Linux, ensuring a broad representation of GUI interactions across both desktop and mobile environments. The application distribution is shown in  Fig. 4(b). Our dataset covers a wide variety of application categories, ensuring a broad and diverse range of user tasks. We show the step distribution in  Fig. 4(c). Our dataset includes GUI interactions with varying step lengths, ranging from single-step actions to 9-step tasks. The distribution shows that shorter tasks (1-3 steps) are more frequent, while longer tasks gradually decrease in numbers. The higher proportion of trajectories of 1-3 steps is partly due to the trajectory splitting strategy described in Section 3.5. The action distribution is shown in  Fig. 4(d). Click and Tap are the most frequent, reflecting their central role in both desktop and mobile interactions. Other common actions include HotKey, Swipe, and Drag, covering various forms of user interactions.
Figure 5: Data flow of data filtering.
We also provide statistics on the data filtering flow in  Fig. 5. Throughout the data filtering pipeline, we observe a gradual reduction in data volume, reflecting the progressive refinement of the dataset. After Duplicate Tutorial Removal, 72.3% of the data remains. Following Content Filtering, this is further reduced to 58.5%. Finally, after applying Trajectory Filtering, 33.0% of the original data is retained. This step-wise filtering process helps to include high-quality and GUI-relevant interactions in the final dataset.
Model Desktop Icon Desktop Text Mobile Icon Mobile Text Web Icon Web Text Average
SeeClick-9.6B [4] 30.0 72.2 52.0 78.0 32.5 55.7 53.4
UGround-7B [7] 63.6 82.5 60.3 82.8 80.4 73.3 70.4
OmniParser-GPT-4V [23] 63.6 91.3 57.0 93.9 51.0 81.3 73.0
ShowUI-2B [18] 61.1 76.3 75.5 92.3 63.6 81.7 75.1
Qwen2.5-VL-3B $\dagger$ 30.7 66.0 40.2 70.0 51.5 69.1 56.5
Qwen2.5-VL-7B $\dagger$ 60.0 86.6 71.6 91.2 67.5 85.2 78.6
TongUI-3B 77.1 92.3 77.7 92.6 74.8 87.8 83.6
TongUI-7B 80.0 93.8 79.5 91.9 81.6 89.1 86.0
@@
Table 1: Results on ScreenSpot; $\dagger$ means the results are re-produced. The best method is marked in bold, and the second-best method is underlined.
Model Mobile Desktop Web Avg
Text Icon/Widget Text Icon/Widget Text Icon/Widget
SeeClick [4] 78.4 50.7 70.1 29.3 55.2 32.5 55.1
OS-Atlas-4B [39] 87.2 59.7 72.7 46.4 85.9 63.1 71.9
OS-Atlas-7B [39] 95.2 75.8 90.7 63.6 90.6 77.3 84.1
Qwen2.5-VL-3B $\dagger$ 77.2 54.5 74.7 35.7 68.8 60.1 64.2
Qwen2.5-VL-7B $\dagger$ 95.5 81.5 90.2 65.0 90.2 70.4 84.0
TongUI-3B 94.4 79.6 92.8 75.0 87.6 77.8 85.5
TongUI-7B 93.1 81.5 96.4 82.9 90.2 84.7 88.7
Table 2: Results on ScreenSpot-V2.
Agent Model Development Creative CAD Scientific Office OS Avg
Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg
GPT-4o 1.3 0.0 0.7 1.0 0.0 0.6 2.0 0.0 1.5 2.1 0.0 1.2 1.1 0.0 0.9 0.0 0.0 0.0 1.3 0.0 0.8
SeeClick [4] 0.6 0.0 0.3 1.0 0.0 0.6 2.5 0.0 1.9 3.5 0.0 2.0 1.1 0.0 0.9 2.8 0.0 1.5 1.8 0.0 1.1
OS-Atlas-4B [39] 7.1 0.0 3.7 3.0 1.4 2.3 2.0 0.0 1.5 9.0 5.5 7.5 5.1 3.8 4.8 5.6 0.0 3.1 5.0 1.7 3.7
ShowUI-2B [18] 16.9 1.4 9.4 9.1 0.0 5.3 2.5 0.0 1.9 13.2 7.3 10.6 15.3 7.5 13.5 10.3 2.2 6.6 10.8 2.6 7.7
CogAgent-18B [10] 14.9 0.7 8.0 9.6 0.0 5.6 7.1 3.1 6.1 22.2 1.8 13.4 13.0 0.0 10.0 5.6 0.0 3.1 12.0 0.8 7.7
Aria-UI [43] 16.2 0.0 8.4 23.7 2.1 14.7 7.6 1.6 6.1 27.1 6.4 18.1 20.3 1.9 16.1 4.7 0.0 2.6 17.1 2.0 11.3
UGround-7B [7] 26.6 2.1 14.7 27.3 2.8 17.0 14.2 1.6 11.1 31.9 2.7 19.3 31.6 11.3 27.0 17.8 0.0 9.7 25.0 2.8 16.5
OS-Atlas-7B [39] 33.1 1.4 17.7 28.8 2.8 17.9 12.2 4.7 10.3 37.5 7.3 24.4 33.9 5.7 27.4 27.1 4.5 16.8 28.1 4.0 18.9
Qwen2.5-VL-3B $\dagger$ 14.9 2.1 8.7 7.1 2.1 5.0 1.5 1.6 1.5 13.9 8.2 11.4 10.7 3.8 9.1 9.4 2.3 6.1 9.1 3.3 6.9
Qwen2.5-VL-7B $\dagger$ 22.1 0.7 11.7 16.2 4.2 11.1 5.1 3.1 4.6 25.0 7.3 17.3 22.0 11.3 19.6 17.8 4.5 11.7 17.4 4.5 12.5
TongUI-3B 32.5 0.7 17.1 24.8 2.8 15.5 11.7 1.6 9.2 43.1 12.7 29.9 32.8 7.6 27.0 15.0 1.1 8.7 26.4 4.1 18.0
TongUI-7B 40.9 3.5 22.7 31.3 7.0 21.1 17.3 9.4 15.3 50.7 12.7 34.3 45.8 13.2 38.3 28.0 6.7 18.4 35.1 8.0 24.7
Table 3: Results on ScreenSpot-Pro.